OpenStudy (anonymous):

Find the slope and y-intercept of 2x-5y=14 ? :)

OpenStudy (anonymous):

please help! (:

OpenStudy (mandre):

Try to make y the subject of the formula. That will give both your answers.

OpenStudy (mandre):

First subtract 2x from both sides. You with me?

OpenStudy (anonymous):

yes :)

OpenStudy (mandre):

What does your equation look like now?

OpenStudy (anonymous):

so y=2/5x+ -14/5......?

OpenStudy (uri):

Formula is y=mx+b So arrange the eq in that way.

OpenStudy (mandre):

so is it -14/5 or +14/5? Can't be both

OpenStudy (anonymous):

-14/5 :)

OpenStudy (mandre):

So now you have y = 2/5x - 14/5 The coefficient of x is the slope and the constant is the y-intercept.

OpenStudy (anonymous):

thanks! :)

OpenStudy (mandre):

pleasure :)

OpenStudy (anonymous):

slope is 2/5 and y-intercept is -14/5